# Awes, Laws, and Flaws From Today’s LLM Research

### v1

@@ Abstract
We perform a critical examination of the scientific methodology behind contemporary large language model (LLM) research. For this we assess over 2,000 research works based on criteria typical of what is considered good research (e.g. presence of statistical tests and reproducibility) and cross-validate it with arguments that are at the centre of controversy (e.g., claims of emergent behaviour, the use of LLMs as evaluators). We find multiple trends, such as declines in claims of emergent behaviour and the presence of ethics disclaimers; and the rise of LLMs as evaluators. This paper underscores the need for more scrutiny and rigour by and from this field. Critical reading and familiarity with the literature are crucial to live up to the fundamentals of a responsible scientific method that is ethical, reproducible, systematic, and open to criticism.
@@ 1 Introduction
It is known that in many areas of AI research there are issues around reproduciblity and other experimentation protocols. For example, research works do not provide sufficient details for independent verification [1, 2]; or they report performances without providing detailed protocols or error breakdowns [3, 1, 4, 5, 2]. This is especially common in health sciences [6], security [7], and recommender systems [8]–all areas where LLMs are increasingly being used. Studies and meta-reviews usually come with suggested good practices like reporting carbon footprints [9]; reproducibility checklists and experimental protocols [10, 11, 12]; self-contained artifacts [13]; and metadata for corpora [14]. These suggestions are often not heeded: Gehrmann et al. [1] recommended good practices for research, and noted that, out of 66 articles from leading conferences, only between 15% and 35% of the recommendations were partially followed. Worse, they practices might not even be sufficiently impactful [7], given that other experimental protocols (e.g., sampling, initialisation, hyperparameters) may also impact reproducibility [3]. This is also relevant to LLMs, as they are intrinsically stochastic and sometimes limited to an API call.
It is then perhaps not surprising that it has been said that the focus of AI research is the approaches, rather than the results–less importance is allocated to issues around experimental protocols when compared to other fields [4]. However, we argue that in LLM research there has been a shift towards result-driven experimentation: partly due to their high availability, but also due to the aforementioned combination of capabilities and attention from stakeholders. This in turn means that the community should allocate the same weight to issues around experimental protocols and good research practices as in other fields.

### v2

@@ Abstract
We perform a critical examination of the scientific methodology behind contemporary large language model (LLM) research. For this we assess over 2,000 research works based on criteria typical of what is considered good research (e.g. presence of statistical tests and reproducibility) and cross-validate it with arguments that are at the centre of controversy (e.g., claims of emergent behaviour, the use of LLMs as evaluators). We find multiple trends, such as declines in claims of emergent behaviour and ethics disclaimers; the rise of LLMs as evaluators in spite of a lack of consensus from the community about their useability; and an increase of claims of LLM reasoning abilities, typically without leveraging human evaluation. This paper underscores the need for more scrutiny and rigour by and from this field to live up to the fundamentals of a responsible scientific method that is ethical, reproducible, systematic, and open to criticism.
@@ 1 Introduction
However, it is known that some areas of AI research have issues with reproduciblity and other experimentation protocols. For example, research works do not provide sufficient details for independent verification [1, 2]; or they report aggregate performances (e.g., accuracy) without providing detailed protocols or error breakdowns [3, 1, 4, 5, 2]. This is especially common in health sciences [6], security [7], and recommender systems [8]–all areas where LLMs are increasingly being used. Even though studies and meta-reviews usually come with suggested good practices (reporting carbon footprints [9]; reproducibility checklists and experimental protocols [10, 11, 12]; self-contained artifacts [13]; and metadata for corpora [14]), these suggestions are often not heeded. Last year, Gehrmann et al. [1] recommended good practices for research, and noted that, out of 66 articles from leading conferences, only between 15% and 35% of the recommendations were partially followed. The practices themselves might not even be sufficiently impactful [7], given that other experimental protocols (e.g., sampling, initialisation, hyperparameters) may also impact reproducibility [3]. These are issues particularly relevant to LLMs, as they are intrinsically stochastic and sometimes limited to an API call.
Given all this, it has been said that the focus of AI research is the approaches, rather than the results, or, rather, that less importance is allocated to issues around experimental protocols when compared to other fields [4]. We argue, however, that in LLM research there has been a shift towards result-driven experimentation: partly due to their high availability and remarkable capabilities, but also due to the aforementioned combination of capabilities and attention from stakeholders. This in turn suggests that the scientific community should allocate the same relevance to experimental protocols and good research practices as in other fields.